OpenStudy (anonymous):

What is the solution to a differential equation which models a simple harmonic movement, with an angular frequency w=(k/m)^(1/2), fase angle=f, and amplitude=A

OpenStudy (anonymous):

Help me out please, answer choices are: a) x = A t sin(w t + f) b) x = A sin(w t + f) c) x = e^(A t) sin(w t + f) d) x = A sin(w t) cos(t - f)